The sum of Annie’s age and Julie’s age is sixty. Julie is four times older than Annie. How old is Julie?
Answer by TimothyLamb(4379)

a + j = 60
j = 4a
---
a + j = 60
a + 4a = 60
5a = 60
---
a = 12
j = 48
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php